OpenStudy (anonymous):

Find the sum to n terms \[\frac{ 1 }{ 1*2}+\frac{ 1 }{ 2*3}+\frac{ 1 }{ 3*4 }+.......\]

OpenStudy (anonymous):

break each term to fractions

OpenStudy (anonymous):

1/2 + 1/6+1/12.......

hartnn (hartnn):

no, write 1 as 2-1 above 2*1 and 3-2 above 3*2.....and so on, then break into fractions.....

OpenStudy (anonymous):

then..)

OpenStudy (anonymous):

@satellite73

hartnn (hartnn):

did u even try?

OpenStudy (anonymous):

Yup...i didnt find any thing..there

hartnn (hartnn):

1-1/2 + 1/2 -1/3 + 1/3 - 1/4 + .... didn't u get this ?

OpenStudy (anonymous):

I did this in another way using kth order difference and got 1/(n+n^2)

OpenStudy (anonymous):

n^2 = n(n+1)(2n+1)/6 n=n(n+1)/2

OpenStudy (shubhamsrg):

you have summation (1/n(n+1) ) now,, 1/n(n+1) = 1/n - 1/(n+1) hope you get it ..

OpenStudy (anonymous):

cant we take it as 1/(n^2+n)

OpenStudy (shubhamsrg):

you can take it any how you want,,but what ever you take ,,you should be comfortable with that,, tell me,,what progress you make with n^2 + n thing?

OpenStudy (anonymous):

n^2 = n(n+1)(2n+1)/6

OpenStudy (anonymous):

n=n(n+1)/2

OpenStudy (shubhamsrg):

you dont have n^2 .. you have 1/n^2 and not only that,,its 1/(n^2 +n) ,,you're confusing yourself,,re-think please..

OpenStudy (anonymous):

Why..

OpenStudy (anonymous):

1/n^2+n = 1/(n(2n+1)(n+1)/6 +n(n+1)/2

OpenStudy (shubhamsrg):

you want to say summation 1/n = 2/n(n+1) ??

OpenStudy (anonymous):

Yes..)why is it wrong

OpenStudy (shubhamsrg):

try this one for n=1 ,2 etc and see for yourself..

OpenStudy (shubhamsrg):

and logically,, summation 1/n = 1/1 + 1/2 + 1/3 ....... and summation n = 1 + 2+ 3+ ...... see it ..is one of them reciprocal of the other ? o.O

OpenStudy (anonymous):

yes..u r correct...)....Continue by ur method

OpenStudy (shubhamsrg):

i already hinted you,,you should try and continue yourself..